# Tian’s theorem for Moishezon spaces

### v1

@@ 1. Introduction
Let $X,\omega,(L_{p},h_{p})$ verify assumptions (A)-(B) and $\pi:\widetilde{X}\to X$ be as in Theorem 2.1. In [CMM] we assumed that $X$ is a Kähler normal space, and we showed that the desingularization $\widetilde{X}$ obtained by finitely many blow-ups with smooth centers as in [BM, GM] is Kähler. This is crucial for construction peak sections by using methods involving $\overline{\partial}$. In our present situation we obtain a projective desingularization $\widetilde{X}$ since $X$ is Moishezon.
@@
This yields the lower bound in (2.2), since $P_{p}(z)\geq|S_{z,p}(z)|^{2}_{h_{p}}/\|S_{z,p}\|_{p}^{2}\geq c^{b_{p}}$. To this end we work first on $\pi^{\star}L_{p}|_{\widetilde{X}\setminus\Sigma}$ using the metric $\widetilde{h}_{p}$ from Lemma 2.3. By (2.1),
@@ 3. Applications
One of the main applications of Tian’s theorem is to the study of the asymptotic distribution of the zeros of random sequences of sections in $H^{0}(Z,L^{p})$ as $p\to\infty$, where $L$ is a holomorphic line bundle over a compact complex manifold $Z$. This started with the pioneering work of Shiffman and Zelditch [SZ1] in the case of a positive line bundle $(L,h)$ over a projective manifold $Z$ (see also [SZ2, Sh]). It is shown in [SZ1] that for almost all sequences $\{\sigma_{p}\in H^{0}(Z,L^{p})\}_{p\geq 1}$ one has that $\frac{1}{p}\,[\sigma_{p}=0]\to c_{1}(L,h)$ weakly on $Z$, where $[s=0]$ denotes the current of integration over the zero divisor of a holomorphic section $s$. In the case of singular Hermitian holomorphic line bundles we proved that similar results hold in different contexts [CM1, CM2, CMN1, CMN2].
@@
We close the paper with some remarks of Moishezon manifolds. By a theorem of Moishezon, a Moishezon manifold is projective if and only if it carries a Kähler metric, see [Mo] and [MM, Theorem 2.2.26]. Moreover, any Moishezon manifold of dimension two is projective, by Theorem 2.1. Indeed, in dimension two we can blow up only points and the blow-up $\widehat{X}$ at a point of a compact manifold $X$ is projective if and only $X$ is projective. Hence non-projective Moishezon manifolds have dimension greater than two. The first example of this kind was obtained by Hironaka in his thesis (1961) and is described in [Ha, Appendix B, Example 3.4.1]. It’s a manifold which contains a curve which is homologous to zero, which is impossible on a Kähler manifold. Further examples can be found in [A, BV, Ko, Pe], see also [MM, Section 2.3.4].

### v2

@@ 1. Introduction
Let $X,\omega,(L_{p},h_{p})$ verify assumptions (A)-(B) and $\pi:\widetilde{X}\to X$ be as in Theorem 2.1. In [CMM] we assumed that $X$ is a normal Kähler space, and we showed that the desingularization $\widetilde{X}$ obtained by finitely many blow-ups with smooth centers as in [BM, GM] is Kähler. This is crucial for construction peak sections by using methods involving $\overline{\partial}$. In our present situation we obtain a projective desingularization $\widetilde{X}$ since $X$ is Moishezon.
@@
This yields the lower bound in (2.2), since $P_{p}(z)\geq|S_{z,p}(z)|^{2}_{h_{p}}/\|S_{z,p}\|_{p}^{2}\geq c^{b_{p}}$. To this end we work first on $\pi^{\star}L_{p}|_{\widetilde{X}\setminus\Sigma}$ using the metric $\widetilde{h}_{p}$ from Lemma 2.3. By estimates (2.1),
@@ 3. Applications
One of the main applications of Tian’s theorem is to the study of the asymptotic distribution of the zeros of random sequences of sections in $H^{0}(Z,L^{p})$ as $p\to\infty$, where $L$ is a holomorphic line bundle over a compact complex manifold $Z$. This started with the pioneering work of Shiffman and Zelditch [SZ1] in the case of a positive line bundle $(L,h)$ over a projective manifold $Z$ (see also [SZ2, Sh]). It is shown in [SZ1] that for almost all sequences $\{\sigma_{p}\in H^{0}(Z,L^{p})\}_{p\geq 1}$, with respect to the spherical measure, one has that $\frac{1}{p}\,[\sigma_{p}=0]\to c_{1}(L,h)$ weakly on $Z$, where $[s=0]$ denotes the current of integration over the zero divisor of a holomorphic section $s$. In the case of singular Hermitian holomorphic line bundles we proved that similar results hold in different contexts [CM1, CM2, CMN1, CMN2].
@@
We close the paper with some remarks of Moishezon manifolds. By a theorem of Moishezon, a Moishezon manifold is projective if and only if it carries a Kähler metric, see [Mo] and [MM, Theorem 2.2.26]. Moreover, any Moishezon manifold of dimension two is projective, by Theorem 2.1. Indeed, in dimension two we can blow up only points and the blow-up $\widehat{X}$ at a point of a compact manifold $X$ is projective if and only if $X$ is projective. Hence non-projective Moishezon manifolds have dimension greater than two. The first example of this kind was obtained by Hironaka in his thesis (1961) and is described in [Ha, Appendix B, Example 3.4.1]. It’s a manifold which contains a curve which is homologous to zero, which is impossible on a Kähler manifold. Further examples can be found in [A, BV, Ko, Pe], see also [MM, Section 2.3.4].